### v2

# Learning Only with Images: Visual Reinforcement Learning with Reasoning, Rendering, and Visual Feedback

Abstract
Multimodal Large Language Models (MLLMs) exhibit impressive performance across various visual tasks. Subsequent investigations into enhancing their visual reasoning abilities have significantly expanded their performance envelope. However, a critical bottleneck in the advancement of MLLMs toward deep visual reasoning is their heavy reliance on curated image-text supervision. To solve this problem, we introduce a novel framework termed “Reasoning-Rendering-Visual-Feedback” (RRVF), which enables MLLMs to learn complex visual reasoning from only raw images. This framework builds on the “Asymmetry of Verification” principle to train MLLMs, i.e., verifying the rendered output against a source image is easier than generating it. We demonstrate that this relative ease provides an ideal reward signal for optimization via Reinforcement Learning (RL) training, reducing reliance on the image-text supervision. Guided by the above principle, RRVF implements a closed-loop iterative process encompassing reasoning, rendering, and visual feedback components, enabling the model to perform self-correction through multi-turn interactions, while this pipeline can be optimized end-to-end by the GRPO algorithm. Extensive evaluations are conducted on image-to-code generation across two diverse domains: data charts and web interfaces. The RRVF-trained model not only outperforms existing open-source MLLMs and supervised fine-tuning baselines but also exhibits superior generalization to unseen datasets. Critically, the model’s performance surpasses that of the more advanced MLLM used to provide the feedback signal during training. This work establishes a self-improvement paradigm that offers a viable path to robust, generalizable models without reliance on explicit supervision. Code will be available at https://github.com/L-O-I/RRVF.
Introduction
Figure 1: Comparison of training paradigms. Conventional MLLMs require paired image-text data for supervised training. The proposed RRVF framework is trained solely on raw images, utilizing a closed-loop process of reasoning, rendering, and visual feedback that is optimized with the GRPO.
Figure 2: An overview of the Reasoning-Rendering-Visual-Feedback (RRVF) framework. The framework is designed to improve visual reasoning using only image inputs. (a) Given an input image, (b) the multimodal language model engages in an iterative reasoning process to generate rendering code. The generated code is executed by external tools to produce a rendered image. (c) The rendered image is compared to the original, and the discrepancies are converted into structured visual feedback. This feedback guides the model in the subsequent reasoning turn. (d) Upon completion of the iterative process, a visual judge quantitatively assesses the similarity between the final rendered output and the original image. (e) This similarity score, along with rewards for format validity and tool usage, is integrated by a reward manager to compute a final reward signal. This signal is used to optimize the model policy via reinforcement learning with the GRPO algorithm.
Improving complex reasoning in Large Language Models (LLMs) represents a fundamental challenge in AI research and a critical milestone on the path to Artificial General Intelligence (AGI) (OpenAI 2024b; DeepSeek-AI et al. 2025; Hu et al. 2025; Team et al. 2025b). Recent works have attempted to replicate this success in Multimodal Large Language Models (MLLMs) for addressing complex visual tasks, yet this paradigm introduces several challenges (Huang et al. 2025; Shen et al. 2025a).
Specifically, for MLLMs, effective reasoning necessitates capabilities that extend beyond elementary visual perception to achieve sophisticated visual comprehension and inference. This paradigm requires models to not only perform object recognition but also to analyze underlying visual semantics, spatial-geometric relationships, and complex inter-object dependencies (Su et al. 2025c; Luo et al. 2024). While accurate visual perception serves as a fundamental prerequisite for such reasoning capabilities, it inherently depends on robust reasoning mechanisms, creating a bidirectional dependency.
Recently, the pioneering work, like OpenAI’s “Think with Images” (OpenAI 2025b; Su et al. 2025a), has shown how external tools like OCR and zoom can enable models to interact proactively with visual content. Subsequent approaches expand beyond limited tool sets by generating executable code for visual operations through Reinforcement Learning (RL) training (Sutton, Barto et al. 1998), as seen in VRAG-RL (Wang et al. 2025b; Liu et al. 2025b). This demonstrates that incorporating external tools and reusing visual representations in MLLM reasoning can effectively enhance model performance.
Inspired by this, we propose an MLLM RL framework called “Reasoning-Rendering-Visual-Feedback” (RRVF) to endow MLLMs with enhanced visual comprehension and reasoning capabilities. As shown in Figure 1, this framework enhances visual understanding through a cycle of reasoning, rendering, and visual feedback, relying solely on images as training data. The core design is motivated by the “Asymmetry of Verification” principle (Wei 2025), i.e., for many tasks, a proposed solution is substantially easier to verify than to generate. Specifically, tasks exhibiting this property are particularly amenable to RL training, as the efficient verification process can be directly formulated as a reward signal. Remarkably, we observe that for image-to-code tasks, e.g., converting the chart image to the executable code, verifying the visual similarity between rendered outputs and source images proves substantially more tractable than generating rendering code from scratch.
Within the RRVF framework, the MLLM reasons about an input image to generate rendering code, which is then executed by a tool. Based on the visual discrepancy between the rendered and original images, the MLLM iteratively refines its code across multiple turns. To optimize this closed-loop refinement process, the system is modeled as a reinforcement learning task driven by the GRPO algorithm (Shao et al. 2024b). The learning is guided by a reward signal that incorporates the final visual similarity score. This mechanism compels the model to understand the image’s generative logic from raw pixels, leading to semantically correct code. The training process requires no text-based ground truth. This naturally resolves the “semantic equivalence” issue, as any program that produces the correct visual output is rewarded.
Moreover, the GRPO optimization is guided by a hybrid reward function. (1) The primary component is a visual similarity reward that quantifies the fidelity between the rendered image and the original image by integrating signals from established vision-language models. (2) This is supplemented by a format correctness reward to ensure the generation of syntactically correct and executable code. And (3) an adaptive tool-use efficiency reward is introduced to balance exploration and convergence. This hybrid reward function effectively enables the model to learn underlying generative logic directly from pixel-level visual representations.
To demonstrate the effectiveness of the proposed RRVF framework, we evaluate our method in two distinct domains, data charts and web interfaces, while using only raw pixel-based images as input. Extensive experiments show that the RRVF framework enables MLLMs to effectively learn the underlying generative logic for both domains, capturing structures that range from the strict syntax of programming toolkits to the flexible layouts of web pages.
Our main contributions are as follows:
  • The exploration of the ”Asymmetry of Verification” principle for MLLM training in complex visual reasoning tasks. We establish that verifying the visual correctness of a rendered output against a source image is a significantly more tractable problem than generating the initial code. This asymmetry provides an effective learning signal that circumvents the need for textual supervision.
  • A novel framework, termed Reasoning-Rendering-Visual-Feedback (RRVF), is proposed. It implements a closed-loop iterative process that enables model self-correction through multi-turn interactions and tool invocation, and this pipeline can be optimized in an end-to-end manner by the GRPO algorithm.
  • Extensive evaluation of the RRVF framework across two structurally diverse domains (i.e., data charts and web interfaces) demonstrates that the proposed approach successfully learns underlying generative logic directly from pixel-level representations. The model trained via RRVF significantly outperforms state-of-the-art open-source MLLMs of comparable scale and surpasses the large-scale model used during training to provide visual feedback and judgment. Notably, it outperforms supervised fine-tuning methods on image-code pairs and exhibits superior generalization capabilities on unseen datasets. This establishes a promising paradigm for developing more capable MLLMs.
Related Works
Multimodal Large Models
Multimodal Large Language Models (MLLMs) are rapidly evolving (Liu et al. 2023; Bai et al. 2025; Zhu et al. 2025; Anthropic 2025; OpenAI 2025a; Google 2025), demonstrating remarkable capabilities in visual understanding and cross-modal interaction. At the forefront are powerful proprietary models such as GPT-o3 (OpenAI 2025a) and Gemini-2.5 (Google 2025), which employ a unified Transformer architecture for end-to-end processing of diverse inputs including images, text, and long videos. Concurrently, the open-source community is making significant strides (Team et al. 2025c; Guo et al. 2025; Team et al. 2025a). For instance, InternVL3 (Zhu et al. 2025) has achieved state-of-the-art (SOTA) performance on numerous multimodal benchmarks through its robust visual encoder and refined alignment mechanism. Similarly, models like Qwen2.5-VL (Bai et al. 2025) have shown strong capabilities in multi-language support and specialized tasks like optical character recognition (OCR) and document understanding. These collective efforts, from both proprietary and open-source domains, have significantly advanced the field and promoted the democratization of related technologies (Chu et al. 2025; Liu et al. 2025a). Despite these advances in visual representation and long-context processing, current MLLMs still have limitations in deep visual reasoning and do not fully utilize visual information (Su et al. 2025c; Wang et al. 2025c; mao2025investigatingredundancymultimodallarge).
Visual Reasoning
Visual reasoning, as a core capability in multimodal understanding, has garnered significant attention from the academic community in recent years (Su et al. 2025c). Efforts to bridge the semantic gap between continuous visual information and discrete language have evolved through several stages: from encoding images into static visual features for Large Language Models (LLMs), to leveraging text-based Chain-of-Thought (CoT) for reasoning (Zhang et al. 2023; Wang et al. 2025c), which explicitly breaks down complex visual problems into textual intermediate steps (Shao et al. 2024a), and more recently, to accomplishing tasks by invoking external tools (Zheng et al. 2025; Xu et al. 2025; Wu et al. 2024b; Fu et al. 2025). This paradigm, where models learn to call APIs, was pioneered in language-only contexts by works like Toolformer (Schick et al. 2023) and extended to multimodal scenarios (Wu et al. 2023). Building upon this, some models will act as advanced planners (Qi et al. 2024; Su et al. 2025b), enriching visual understanding at a finer-grained level by dynamically and iteratively invoking tools like ”zoom in” and ”crop” as reasoning steps. To further enhance the model’s autonomous cognitive abilities, subsequent research has transcended the limitations of predefined toolsets by generating executable visual code (Beltramelli 2018; Zhao et al. 2025). This approach increases reasoning flexibility by decomposing complex visual tasks into modular and programmatic sub-tasks (Wang et al. 2025a; Mallis et al. 2024; Hu et al. 2024; Shen et al. 2025b; Sur´ıs, Menon, and Vondrick 2023). Despite these advancements, robust and general-purpose multimodal reasoning remains an open challenge, with substantial room for future exploration.
Method
This section details the Reasoning-Rendering-Visual-Feedback (RRVF) framework first, then the reinforcement learning strategy for optimization.
RRVF Framework
The core of RRVF is a closed-loop system comprising three key components: an iterative visual reasoner, a visual feedback mechanism, and a final visual judge.
Algorithm 1 Iterative Visual Reasoning
1:MLLM $\mathcal{M}$, Single Image $I$, Max Turns $T_{max}$
2:Final rendering code $C_{final}$
[…]
4:for $t=1$ to $T_{max}$ do
5:  $R_{model}\leftarrow\mathcal{M}.\text{Generate}(H)$triangleright Generate response
6:  $H\leftarrow H\oplus R_{model}$
7:  if $R_{model}$ contains <answer> then
[…]
9:   break $\triangleright$ Final turn: task is complete
10:  else
[…]
12:   $F\leftarrow\text{CallTool}(C_{tool})$triangleright Invoke the external tool
13:   $H\leftarrow H\oplus F$triangleright Append visual feedback to history
14:  end if
15:end for
16:return $C_{final}$
Visual Reasoning.
The process for generating rendering code is an iterative procedure, formally outlined in Algorithm 1. It begins with a single image input. In each turn, as illustrated in Figure 2(b), the MLLM produces a response (Line 3) that includes both internal reasoning in <think> tags and a specific action. For intermediate turns, this action is a code snippet enclosed in a <tool_call> tag. A call is then made to an external tool to execute this code (Line 10). The resulting feedback is appended to the conversation history to inform the next turn. This iterative cycle concludes when the model’s response contains the final solution within an <answer> tag, which terminates the process (Lines 5-7). The process also stops if a predefined maximum number of turns is reached (Yao et al. 2023; Madaan et al. 2023).
Visual Feedback.
The visual feedback mechanism (Bai et al. 2022), illustrated in Figure 2(c), is a critical tool that provides structured guidance for the MLLM. Its operation involves two main steps:
  • Step 1: Rendering. The code generated by the MLLM is executed by a specific rendering engine based on the domain. For data charts, Python code is executed using libraries such as Matplotlib. For web interfaces, the Playwright library is utilized to render the HTML in a browser and capture a screenshot. In cases where the code is invalid and rendering fails, the tool returns an explicit failure signal.
  • Step 2: Comparison and Feedback Generation. The successfully rendered image is compared to the original input image. This comparison is performed by a separate, powerful MLLM acting as a qualitative assessor. It is prompted to articulate the specific visual discrepancies (e.g., colors, elements, missing text) in natural language. This descriptive feedback provides actionable information for the primary MLLM’s next reasoning step.
Visual Judge.
The Visual Judge (Figure 2(d)) provides a quantitative score for the final output, which is essential for policy optimization. The process first renders the final code and then uses a powerful MLLM to evaluate the similarity between the generated and target images (Zheng et al. 2023). The MLLM-based score serves as one of the reward signals for reinforcement learning and forms the primary optimization objective.
Table 1: Performance comparison on the ChartMimic benchmark. We report compilation success (“Exec rate”), similarity scores for specific attributes, and a quality score from “GPT-4o”, which are the metrics from the original ChartMimic benchmark (Yang et al. 2024). The best and second-best results among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
Model Exec rate Text Layout Type Color GPT-4o score Overall
Closed-Source MLLMs
(2024/02) Gemini-1.0-Pro-Vision 68.2* 52.6* 64.2* 51.3* 47.1* 53.3* 53.6*
(2024/11) GPT-4o-2024-11-20 90.00 66.55 79.31 71.83 60.84 82.50 76.06
(2025/04) OpenAI O3 90.17 74.17 80.58 71.37 63.74 86.45 79.46
(2025/05) Claude-4-Sonnet 91.83 68.87 82.43 67.13 57.59 85.46 77.23
(2025/06) Gemini-2.5-Pro 93.33 84.95 83.37 75.05 66.90 90.58 84.07
Open-Source MLLMs
(2025/02) Qwen2.5-VL-72B-Instruct 83.83 34.44 61.71 45.49 35.12 50.41 47.30
(2024/03) DeepSeek-VL-7B 41.3* 15.3* 26.6* 19.7* 14.5* 20.4* 19.7*
(2025/02) LLaVA-OneVision-7B 17.28 7.97 13.55 9.15 7.36 10.01 9.76
(2025/02) Qwen2.5-VL-7B-Instruct 68.83 30.01 55.79 36.50 26.91 39.04 38.17
(2025/04) InternVL3-8B 71.67 45.03 57.89 45.87 38.88 54.91 50.91
SFT [with text labels] 69.00 56.97 63.60 60.53 51.89 62.09 60.17
$\Delta$ (vs Qwen2.5-VL-7B-Instruct) +0.17 +26.96 +7.81 +24.03 +24.98 +23.05 +22.00
RRVF (Ours) [without text labels] 97.83 62.47 80.97 53.56 46.41 67.87 64.36
$\Delta$ (vs Qwen2.5-VL-7B-Instruct) +29.00 +32.46 +25.18 +17.06 +19.50 +28.83 +26.19
Table 2: Performance comparison on the Plot2Code benchmark. The best and second-best results on the primary metrics (Exec Rate, Text, GPT-4o Score) among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
[…]
Closed-Source MLLMs
(2023/09) GPT-4V 84.1* 48.53* 5.45* 57.7* 6.48*
(2024/02) Gemini-1.0-Pro-Vision 68.2* 36.56* 3.45* 53.6* 5.06*
(2024/06) Claude-3-Sonnet 75.8* 35.40* 4.08* 46.7* 5.38*
(2024/11) GPT-4o-2024-11-20 90.15 48.91 6.09 54.25 6.76
(2025/04) OpenAI O3 87.12 57.65 6.70 66.17 7.69
(2025/05) Claude-4-Sonnet 92.42 56.86 6.16 61.52 6.76
(2025/06) Gemini-2.5-Pro 87.88 71.70 7.65 81.59 8.71
Open-Source MLLMs
(2025/02) Qwen2.5-VL-72B-Instruct 83.33 56.74 5.79 68.09 6.95
(2024/03) Mini-Gemini-8x7B-HD 73.5* 29.91* 2.84* 40.7* 3.87*
(2025/02) LLaVA-OneVision-7B 84.09 26.72 2.75 31.78 3.27
(2025/02) Qwen2.5-VL-7B-Instruct 70.46 35.80 3.40 50.81 4.82
(2025/04) InternVL3-8B 76.52 30.67 3.25 40.08 4.25
SFT [with text labels, ChartMimic trained] 49.24 21.63 2.47 43.93 5.02
$\Delta$ (vs Qwen2.5-VL-7B-Instruct) -21.22 -14.17 -0.93 - -
RRVF (Ours) [without text labels] 96.21 39.89 4.44 41.46 4.61
$\Delta$ (vs Qwen2.5-VL-7B-Instruct) +25.75 +4.09 +1.04 - -
Table 3: Performance comparison on the WebSight benchmark for web interface generation. The best results among open-source models under 10B parameters are bolded.
Model CLIP Score GPT Score
Closed-Source MLLMs
GPT-4o-2024-11-20 88.94 94.55
OpenAI O3 91.58 96.49
Claude-4-Sonnet 92.30 96.46
Gemini-2.5-Pro 77.83 75.88
Open-Source MLLMs
LLaVA-OneVision-7B 79.74 72.61
Qwen2.5-VL-7B-Instruct 83.50 84.17
InternVL3-8B 84.17 85.54
RRVF (Ours) 88.29 91.50
Reinforcement Learning Optimization
For image-to-code generation tasks, the iterative and goal-oriented characteristics of the proposed RRVF framework render it particularly amenable to reinforcement learning formulations (Ouyang et al. 2022; Chen et al. 2021).
The optimization process can be effectively formulated as learning an optimal policy that generates accurate rendering code, thereby enhancing the model’s fundamental visual understanding capabilities. Additionally, Group Relative Policy Optimization (GRPO) (Shao et al. 2024b) has been demonstrated to effectively guide model learning toward optimal trajectories through objective verification. Therefore, we employ GRPO as the optimization algorithm for RRVF to learn the capability of modeling rendering code by verifying the visual similarity between rendered outputs and source images. In the following sections, we provide a detailed description of our optimization algorithm.
Optimization with GRPO.
The MLLM’s policy, denoted as $\pi_{\theta}$, is optimized using GRPO, which serves as a computationally efficient alternative to Proximal Policy Optimization (PPO) (Schulman et al. 2017). Unlike PPO, GRPO eliminates the need for a separate value function for advantage estimation, thereby avoiding high computational overhead and streamlining the optimization process by generating a set of candidate outputs from the previous policy iteration.
Specifically, GRPO samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ for each input query $q$ from the old policy $\pi_{\theta_{\mathrm{old}}}$. The method then optimizes the policy by maximizing a surrogate objective that incorporates clipping for stability, along with a KL regularization term to prevent excessive deviation from a reference policy $\pi_{\text{ref}}$. The objective function is given as:
$J_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(O\mid q)}$ (1)
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\!\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)}\,A_{i},\,$
[…]
$-\,\beta\,D_{\mathrm{KL}}\bigl{(}\pi_{\theta}\,\big{\|}\,\pi_{\mathrm{ref}}\bigr{)}\Biggr{]},$
The advantage $A_{i}$ is derived as a normalized relative measure: $A_{i}=\frac{r_{i}-\mathrm{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\mathrm{std}(\{r_{1},r_{2},\dots,r_{G}\})}$, using the rewards from the sampled group to establish a baseline. Additionally, the KL divergence term $D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$ promotes controlled policy updates, balancing exploration and stability throughout training.
Hybrid Reward Design.
As depicted in Figure 2(e), a hybrid reward function is engineered to provide a comprehensive learning signal. Our reward design necessitates the provision of dense and accurate reward signals to ensure format compliance during the model’s inference process while encouraging tool utilization to enhance the model’s performance ceiling. To address these objectives, we design the following three reward components:
  • Visual Similarity Reward ($R_{\text{vision}}$): This is the primary reward component, sourced directly from the Visual Judge module. It quantifies the fidelity between the final rendered image and the original input, providing a dense and accurate signal of task success. This reward can be efficiently computed using an MLLM as a judge, thereby avoiding heavy reliance on curated image-text supervision.
  • Format Correctness Reward ($R_{\text{format}}$): This component provides a crucial binary signal that encourages the generation of valid and executable code. A positive reward is granted if the MLLM’s final code is syntactically correct and can be executed by the rendering engine without errors. Otherwise, a penalty is applied. This reward discourages the model from generating nonsensical or incomplete programs.
  • Tool-Use Reward ($R_{\text{tool}}$): This reward is designed to encourage tool use. It provides a small reward for each successful tool call, incentivizing the model to use the feedback loop to refine its answer. To avoid rewarding excessively long conversations, the reward is capped, balancing the need for exploration with the goal of convergence.
Furthermore, during RRVF training, we employ a weighted combination to balance these reward functions. The total reward $R$ for a completed trajectory is a weighted sum of three distinct components:
$R=w_{v}R_{\text{vision}}+w_{f}R_{\text{format}}+w_{t}R_{\text{tool}},$ (2)
where $w_{v},w_{f}$, and $w_{t}$ are hyperparameters that balance the contribution of each component.
Experiments
Datasets and Evaluation
Datasets and Metrics.
The framework’s performance is evaluated on two image-to-code tasks: chart-to-code and web-to-code. For the chart-to-code task, two standard benchmarks are utilized. The official “chart-to-code” subset of the ChartMimic dataset (Yang et al. 2024) comprises 2,400 test images. From this set, 1,800 images are repurposed for training, while the remaining 600 images, which form the official smaller test set, are used for evaluation. Additionally, the Plot2Code dataset (Wu et al. 2024a), comprising 368 test cases, is used for zero-shot evaluation. For the web-to-code task, only 1,000 screenshots are randomly sampled from the WebSight dataset (Laurençon, Tronchon, and Sanh 2024) for training, with a disjoint set of 500 images reserved for testing.
For the chart-to-code task, evaluation adheres to the official protocols and metrics of its respective benchmarks. For the web-to-code task, however, traditional metrics such as block-level HTML comparisons are ill-suited for our framework. This is because the model is designed to generate semantically equivalent but structurally distinct code, without access to the ground-truth source. Consequently, evaluation for this task is based on visual fidelity and semantic correctness, quantified using the following metrics. (1) CLIP Similarity: The perceptual similarity between the rendered image and the original input is measured by the cosine similarity of their respective CLIP embeddings (Radford et al. 2021). (2) LLM-based Assessment: A powerful multi-modal model, GPT-4o (OpenAI 2024a), is employed as an impartial judge to assess the quality of the generated outputs. It evaluates semantic consistency, element completeness, and overall visual fidelity against the input image.
A comparison with conventional supervised fine-tuning approaches is presented in the ablation study.
Baselines.
A comprehensive set of both closed-source and open-source models is selected for comparison to ensure a thorough evaluation. The closed-source baselines include leading proprietary models: GPT-4o (OpenAI 2024a), OpenAI O3 (OpenAI 2025a), Gemini-1.0-Pro-Vision (Google 2024), Gemini-2.5-Pro (Google 2025), and Claude-4-Sonnet (Anthropic 2025). The open-source counterparts include vision-language models such as Qwen2.5-VL (Bai et al. 2025), LLaVA-OneVision(Li et al. 2024a), Mini-Gemini (Li et al. 2024b), DeepSeek-VL (Lu et al. 2024), and InternVL3 (Zhu et al. 2025).
Experimental Settings
Training Setups.
The policy model is initialized from Qwen2.5-VL-Instruct-7B (Bai et al. 2025) and trained on a cluster of 8 NVIDIA A100 GPUs. The model is optimized using the GRPO algorithm with a global batch size of 32 and a learning rate of $1\times 10^{-6}$. During each optimization step, a group of $G=8$ candidate trajectories is sampled for each prompt. The RRVF interaction loop is set to a maximum of 4 turns, and the maximum sequence length for generation is 16,384 tokens. The KL divergence coefficient is set to 0.0.
Visual feedback and judgment are provided by a more capable model, Qwen2.5-VL-Instruct-72B (Bai et al. 2025), which serves as the reward oracle. To handle the substantial computational demand during the policy rollout phase, this judge model is deployed on a separate, dedicated cluster of 8 NVIDIA A100 GPUs. Inference is accelerated using the vLLM framework (Kwon et al. 2023) to ensure high throughput for the large volume of verification requests generated in this process.
The optimization is guided by the hybrid reward function from Equation 2. The weights are set to $w_{v}=-0.2$, $w_{f}=0.8$, and $w_{t}=1.0$. The interaction loop operates for a maximum of $T_{max}=4$ turns. Within this loop, the tool-use efficiency reward, $R_{\text{eff}}$, assigns a reward of 1/3 for each successful tool call and is fixed to its maximum value of 1.0 if the visual similarity score surpasses 0.95.
Inference Setups.
During the evaluation phase, the model operates in a direct, single-turn inference mode. The final output is generated from the input prompt in a single pass, without access to the iterative refinement mechanism or any external tools available during the training loop. This protocol is intentionally designed to rigorously assess the model’s internalized multi-modal comprehension and generation capabilities, demonstrating the improvements gained autonomously after the completion of training.
Main Results
The main experimental results are presented in Table 1, Table 2, and Table 3. The findings indicate that the RRVF-trained model achieves state-of-the-art performance among open-source Multimodal Large Language Models (MLLMs) of comparable size. This consistent outperformance across diverse tasks highlights the efficacy and generalizability of the proposed framework, which learns complex visual-to-code logic solely from visual feedback.
Chart-to-Code Task.
On the ChartMimic benchmark (Table 1), the RRVF-trained model exhibits substantial improvements over its base model, Qwen2.5-VL-7B-Instruct. The most notable result is the code execution rate (Exec rate) of 97.83%, which surpasses all other evaluated models, including proprietary systems like OpenAI O3 and Gemini-2.5-Pro. This exceptional reliability in generating valid code is a direct result of the RL process, which penalizes non-executable outputs. Consequently, the model attains an Overall score of 64.36, securing the highest performance among all open-source models. This performance is further supported by a strong score in layout understanding (Layout: 80.97), suggesting that the visual feedback loop effectively teaches the model to infer and replicate the underlying structural logic of data visualizations.
On the Plot2Code dataset (Table 2), the RRVF model achieves the highest execution rate (96.21%) among all open-source competitors. This outstanding reliability is directly reflected in the primary metrics (Text and GPT-4o Score), which penalize execution failures by being multiplied by the execution rate. On these realistic metrics, RRVF surpasses all similar-sized models, underscoring its enhanced robustness and reliability. When considering “pass” scores, which are calculated only on successfully executed code, models with lower execution rates are effectively evaluated on easier subsets of data. Despite RRVF’s high execution rate exposing it to more challenging examples, its performance remains highly competitive.
Web-to-Code Task.
The applicability of RRVF is further tested on the WebSight benchmark, a domain characterized by high structural freedom (Table 3). In this task, our RRVF-trained model achieves strong performance with a CLIP score of 88.29 and a GPT score of 91.50, surpassing the similar-sized models. Crucially, this advantage is obtained without any access to ground-truth HTML, relying instead on our proposed method. This demonstrates that the framework can effectively learn to deconstruct and replicate complex layouts and component relationships from raw pixels alone, validating the potential of learning through pure visual feedback even in highly unstructured domains.
Ablation Study
RRVF vs. SFT.
Our ablation study, conducted on the ChartMimic dataset, directly pits RRVF against a standard Supervised Fine-Tuning (SFT) baseline. Although SFT benefits from direct access to ground-truth code, RRVF demonstrates marked superiority. As shown in Table 1, RRVF achieves a near-perfect execution rate of 97.83%, massively outperforming SFT’s 69.00%. This highlights the reliability of the code generated by our framework. Furthermore, RRVF’s dominant Layout score (80.97 vs. 63.60) confirms its advanced comprehension of holistic visual structures. This indicates that learning from the final visual outcome, as RRVF does, cultivates a more robust and compositional reasoning ability compared to the surface-level mimicry of SFT. A more profound finding is that our final 7B model, trained via RRVF, surpasses its much larger 72B teacher model in the Overall score (64.36 vs. 47.30). This counter-intuitive result proves that RRVF is more than knowledge distillation; it enables the policy model to actively explore the solution space and discover policies that are superior to its own feedback provider, thus bootstrapping its capabilities beyond initial constraints.
Generalization Ability.
To evaluate the true extent of the learned skills, we test the models’ generalization ability via a zero-shot evaluation on the unseen Plot2Code dataset. The SFT-trained model exhibits a classic case of overfitting; its execution rate dramatically plummets from 69.00% on the training domain to a mere 49.24% on the new dataset. This sharp decline suggests that SFT learns brittle, non-transferable knowledge tied to the specific code patterns of the ChartMimic dataset. In stark contrast, our RRVF-trained model showcases exceptional generalization. Its execution rate remains remarkably stable, decreasing only negligibly from 97.83% to 96.21%. This stability is strong evidence that RRVF learns the fundamental and transferable principles of translating visual elements into programmatic logic, rather than simply memorizing code templates. This result underscores a key advantage of our framework: it fosters robust reasoning skills that reliably generalize to new and unfamiliar challenges.
Limitations
The framework’s applicability is currently limited to code reconstruction tasks. Investigating the adaptation of RRVF to broader tasks, particularly those involving abstract reasoning where verification is non-trivial, remains an important avenue for future work.
Conclusions
This paper presents RRVF, a framework that enhances visual reasoning in multimodal large language models through a reasoning-rendering-visual-feedback loop. The framework leverages the ”Asymmetry of Verification” principle to learn generative logic directly from raw images, eliminating the need for paired textual supervision. Comprehensive experiments across charts and web interfaces show that RRVF consistently outperforms strong open-source baselines. The ablation study further confirms the superiority of the proposed framework over traditional supervised fine-tuning, validating its effectiveness in fostering robust visual reasoning.